# Augmenting LLMs for General Time Series Understanding and Prediction

### v1

@@ 4.1 Evaluations
Table 3: Time series classification accuracy on the UCR-TSC benchmark collection. TsLLM achieves competitive performance (81.2% mean accuracy with few-shot prompting) without task-specific architectures, demonstrating effective transfer of language model capabilities to temporal pattern recognition.
Metric DTW CNN T-Loss TS2Vec GPT4TS Times- MOM- TsLLM TsLLM
@@
Mean 0.764 0.751 0.833 0.851 0.566 0.572 0.794 0.792 0.812
Median 0.768 0.773 0.849 0.871 0.583 0.565 0.815 0.798 0.823
Forecasting
We evaluate forecasting performance on the GIFT-Eval dataset. Performance is measured using a relative Mean Absolute Percent Error (rMAPE) normalized against the seasonal naive baseline. All baseline results are as reported in Aksu et al. (2024). We further break down performance by the number of variates (univariate vs multivariate) and prediction length, as illustrated in Table 4, where TsLLM performs competitively with other models although not completely outperforming all.
Table 4: Forecasting performance on GIFT-Eval benchmark across different data characteristics. Values show relative MAPE (lower is better) normalized to seasonal naive baseline, following (Aksu et al., 2024). TsLLM with few-shot prompting (FS) achieves best-in-class performance on univariate series (0.78) while remaining competitive overall (0.87), with particularly strong results on long-horizon forecasts.
@@
4.2 Ablation Studies
Table 5: Forecast performance on GIFT-Eval for each experiment in our ablation study, as measured by relative MAPE.
Ablation Value rMAPE Relative Change
Baseline 0.94
LLM SmolLM2-135M 0.99 -5.5%
LLM SmolLM2-1.7B 0.92 +2.1%
LLM Llama 3.2 3B 0.89 +5.3%
Adapter Linear 0.94 -0.2%
Adapter Cross Attention 0.96 -2.4%
Encoder VQ-VAE 1.24 -31.9%
In order to analyze the performance characteristics of different aspects of the TsLLM architecture and training procedure, we conduct ablation studies on our methodology at a smaller scale. In Table 5, we report the results of our primary findings. The baseline for this experiment uses SmolLM2-360M as the pretrained LLM backbone with an MLP-based adapter for the time series encoder and decoder models, and we utilize the same encoder-decoder pair as in previous experiments. These results demonstrate that the size and performance of the backbone LLM matter significantly, suggesting that TsLLM can be scaled up further to improve performance. They also show that a single linear layer performs nearly as well as the MLP adapter, although we have seen the MLP improves performance on other tasks. Finally, using an encoder-decoder pair with quantized latent embeddings currently has a big negative impact on performance, although they are commonly used in vision-language models. This likely could be improved in future work, but highlights the importance of precise, high-fidelity encoder and decoder models for time series data.
Time Series Encoder-Decoder Training
An ablation study on time series encoder-decoder reconstruction performance (Table 6) reveals a critical trade-off between the reconstruction performance of the encoder-decoder pair and downstream performance on practical tasks. We look at the reconstruction MSE on our VAE validation data across encoder-decoder pairs with different hyperparameter settings, as well as the accuracy of small-scale implementations of TsLLM trained using each of these pairs. This table makes clear that it is possible to improve the reconstruction ability of the VAE significantly by reducing the patch size or $\beta$, but this leads to more granular representations of the time series that are not as informative or effective when integrated into TsLLM. The representation abstraction can also be increased by increasing $\beta$ or selecting a larger patch size, but this hurts reconstruction, which also hurts performance of TsLLM. We have optimized for a balance between these objectives in order to maximize performance of TsLLM on real-world tasks.
Table 6: Trade-off between reconstruction quality and downstream task performance for different encoder-decoder configurations. Lower reconstruction MSE does not guarantee better ECG-QA accuracy – our $\beta$-VAE baseline achieves optimal task performance despite higher reconstruction error than a standard autoencoder. Color coding: green indicates improvement, red indicates degradation relative to baseline.
Ablation Baseline Patch size $\beta$ Objective # of layers
Value 16 64 1.0 1.8 AE VQ-VAE 12 48
Recon. MSE 0.217 0.151 0.308 0.173 0.367 0.005 0.432 0.217 0.216
% Change -30% +42% -20% +69% -97% 99% 0% -1%
ECG-QA Acc. 55.9 54.1 53.5 55.5 55.3 54.0 51.8 55.6 54.9
% Change -3.2% -4.3% -0.7% -1.1% -3.4% -7.3% -0.5% -1.8%
5 Discussion
Training LLMs to understand and analyze numerical time series data represents a paradigm shift in time series analysis, moving beyond traditional approaches that treat numerical data in isolation. By augmenting an LLM with time series perception capabilities, we enable new analytical workflows at the intersection of temporal reasoning and natural language understanding. This approach parallels the successful development of vision-language models, which have proven valuable not by achieving state-of-the-art performance on traditional computer vision benchmarks, but by enabling entirely new capabilities such as visual question answering, image captioning, and multimodal reasoning. Similarly, TsLLM’s value lies not in marginal improvements on forecasting accuracy or classification F1 scores, but in its ability to answer questions about temporal patterns, incorporate unstructured contextual information into predictions, and provide natural language explanations for time series phenomena.
The implications of this paradigm shift extend beyond technical capabilities. TsLLM democratizes time series analysis by allowing domain experts – clinicians, financial analysts, environmental scientists – to directly query and understand temporal data through natural language, without requiring specialized data science training. A cardiologist can ask “What abnormalities are present in this ECG, and are they consistent with the patient’s medication history?” A supply chain analyst can request “Forecast next month’s demand considering recent customer reviews and competitor actions.” These interactions were previously impossible with traditional time series models, which require carefully structured numerical inputs and produce only numerical outputs.
Our approach is fundamentally complementary to, rather than competitive with, specialized time series models. Just as vision-language models have not replaced CNNs for image classification but have opened new application domains, TsLLM does not aim to replace ARIMA models or neural forecasters for pure numerical prediction tasks. Instead, it addresses a different class of problems: those requiring the integration of temporal analysis with contextual understanding, domain knowledge, and natural language interaction. This positioning explains why we do not emphasize traditional benchmark performance as our primary evaluation metric – excelling at these benchmarks is not our goal, just as vision-language models do not target ImageNet state-of-the-art as their primary objective.
Limitations
Several limitations of TsLLM warrant discussion. First, TsLLM requires substantially more computational resources than traditional time series models, both for training and inference. This is an expected tradeoff when leveraging the capabilities of large language models – similar to how vision-language models require vastly more training than specialized image classifiers. The benefit is that TsLLM can be trained once and applied to diverse tasks through in-context learning, amortizing costs across applications and eliminating the need for task-specific model development. Second, while TsLLM demonstrates competitive performance on traditional benchmarks, it does not achieve state-of-the-art results on pure numerical tasks like forecasting or anomaly detection. As previous work has noted (Tan et al., 2024; Merrill et al., 2024), transformers, and LLMs in particular, are not ideally suited for numerical time series data, and we have not overcome this limitation. However, our goal is not to compete with specialized models optimized for specific metrics, but to enable capabilities that those models cannot provide. The performance gap on traditional benchmarks is acceptable given TsLLM’s unique ability to incorporate contextual information, go beyond the standard time series tasks, and answer questions. Our current approach to multivariate time series also has limitations. While effective for independent signals, the model may not fully capture complex interrelationships among variables. This reflects our design choice to handle multivariate complexity within the LLM rather than the encoder, prioritizing flexibility over specialized architectural optimizations. Data availability remains a challenge. High-quality datasets combining time series with natural language descriptions are scarce, requiring us to create our own collection through synthesis and curation. While we have assembled the largest publicly available dataset of interleaved time series and text, it represents only a fraction of potential applications. Expanding coverage to new domains requires substantial effort. Finally, evaluation methodology for time series-language tasks lacks standardization. Unlike pure time series or pure language tasks, there are no established benchmarks for assessing models that bridge both modalities. Existing attempts often rely on synthetic data with questionable quality. Developing rigorous, expert-validated benchmarks for tasks like time series question-answering and contextual forecasting should be a priority for the field.
6 Conclusion
This work presents TsLLM, a large language model augmented with the ability to understand and generate numerical time series signals. By introducing a lightweight patch-based encoder-decoder pair and adapter layers, we enable a pretrained LLM to process interleaved numerical signals and text without causing major degradation to its existing capabilities. A unified autoregressive training objective allows TsLLM to tackle forecasting, classification, anomaly detection and question answering in a single framework, and extensive experiments demonstrate competitive performance on traditional numerical benchmarks while unlocking tasks that require joint temporal and linguistic reasoning.
Our findings highlight three main take-aways. First, high-fidelity numerical encodings are essential: representations that excel at reconstruction translate directly into better downstream performance. Second, the quality of the backbone LLM remains a dominant factor; stronger language models amplify the benefits of specialised perception modules. Third, few-shot prompting can be a practical substitute for task-specific fine-tuning, suggesting that future work can focus on data quality rather than ever larger supervised corpora.
Despite these advances, TsLLM is not a replacement for purpose-built time series models. It lags behind state-of-the-art forecasting models on long-horizon multivariate benchmarks and is much more computationally expensive. We therefore view TsLLM as complementary: it excels when time series data must be interpreted in the context of domain knowledge or when natural-language interaction is required.
Looking ahead, we believe the most fruitful directions lie in (i) better handling of multivariate dependencies, (ii) tighter integration of structured domain knowledge, and (iii) community benchmarks that evaluate combined numerical-linguistic competence. We release code, models and data to facilitate such work and hope that TsLLM serves as a stepping stone toward more capable, transparent and deployable time series foundation models.
Appendix A Implementation Details
We implemented TsLLM using PyTorch and Transformers, with FlashAttention-2 and fully sharded data parallelism for efficient distributed training. The complete training process consumed approximately 30 billion tokens across all stages, requiring approximately 1000 GPU-hours on NVIDIA H100s. Our primary experiments use Qwen2.5-7B [Wang et al., 2024] as the base LLM, chosen for its strong performance on standard benchmarks. We also validated our approach with smaller models and different architectures (SmolLM2-360M [Allal et al., 2025], Llama-3.2-3B [Grattafiori et al., 2024]), before scaling up to this size.
The encoder employed a patch-transformer architecture with 24 layers, an embedding size of 64, and a patch size of 32. Time series inputs were standardized using Z-score normalization. The VAE’s latent space is 16-dimensional, and the KL divergence loss was scaled with a peak $\beta$ coefficient of 1.5. VAE training was conducted for 100 million samples using a batch size of 2048. The optimizer was AdamW, with a learning rate of 5e-4 and a linear warmup scheduler. The mean squared error (MSE) loss was used for reconstruction.
TsLLM was optimized using AdamW, with a learning rate of 1e-5 and gradient norm clipping set to 1.0. The total training dataset comprised 20 billion tokens, with a sequence length of 4096 and a micro-batch size of 4. MSE loss was used for time series next patch prediction on time series tokens, and cross entropy loss was used on all non-padding tokens.
Appendix B Additional Results
@@
We also evaluate TsLLM on anomaly detection, specifically on the MSL dataset [Hundman et al., 2018], as shown in Table 7. While our approach achieves a competitive F1 score of 73.3, there remains a performance gap compared to state-of-the-art models like UniTS-FT (81.2) and iTransformer-FT (80.4). This suggests that while TsLLM’s general-purpose architecture provides reasonable anomaly detection capabilities, there may be room for incorporating specialized mechanisms to better identify temporal anomalies. The results demonstrate TsLLM’s versatility across different time series tasks, though specialized models may still hold advantages in specific domains.
@@ Appendix C Datasets
C.1 Training Data Overview
Table 8 summarizes the key statistics for all datasets used in training TsLLM. Our training corpus encompasses over 8 billion total tokens across 20 distinct datasets, spanning diverse domains including medical signals, financial data, environmental measurements, and industrial sensors. The datasets support multiple task types including forecasting, classification, anomaly detection, segmentation, and question-answering, enabling TsLLM to develop comprehensive time series analysis capabilities.
@@ C.1 Training Data Overview
ECG-QA ECG Q-A 110,278,285 607,619 (34,289,034/75,381,632) 159,306
GIFT Multi-Domain Forecasting 10,380,569 7,724,958 (0/2,655,611) 96,307
LOTSA Multi-Domain Forecasting 7,951,924,108 5,961,122,236 (0/1,990,801,872) 74,805,850
@@
PTBXL-Classific. ECG Classification 20,615,930 296,710 (3,577,588/16,741,632) 21,799
PTBXL-Forecast. ECG Forecasting 22,482,365 5,740,733 (0/16,741,632) 21,799
PTBXL-Rep. Gen. ECG Report Gen. 23,155,582 3,689,075 (2,724,875/16,741,632) 21,799
GIFT-Eval (Val) Multi-Domain Forecasting 39,338,366 27,820,967 (0/11,517,399) 346,955
@@
Synthetic (Ours) Multi-Domain Mixed 7,481,857,697 7,011,433,500 (368,202,956/102,221,241) 6,273,306
@@
C.2 Unimodal Time Series Data
@@
For large-scale pretraining, we utilize the GIFT-Eval benchmark [Aksu et al., 2024], a comprehensive collection of 144,000 time series across 23 datasets spanning seven domains. GIFT-Eval provides dedicated splits for pretraining, training, and testing, making it particularly suitable for developing and evaluating general-purpose time series models. The benchmark’s temporal granularity ranges from secondly to yearly frequencies, incorporating both univariate and multivariate sequences with forecast horizons varying from short-term to long-term predictions. With over 230 billion data points, this dataset provides the scale and diversity necessary for training foundation models. We use it extensively for pre-training our time series encoder and decoder models, as well as training TsLLM for forecasting tasks, while carefully ensuring no data leakage into the GIFT-Eval test set used in our evaluation (Section 4).
@@
For time series classification tasks, we employ two principal benchmark collections that have become standard in the field. The UCR Time Series Classification Archive [Dau et al., 2018] comprises 128 datasets spanning diverse applications including electrocardiograms, motion capture sequences, sensor readings, and spectrographs. These datasets vary significantly in length, class distribution, and domain-specific characteristics, providing comprehensive coverage for univariate time series classification. Complementing this, the UEA Time Series Classification Archive [Bagnall et al., 2018] extends our capabilities to multivariate time series, offering 30 datasets that capture complex temporal patterns in applications such as human activity recognition, motion capture, and handwriting analysis. Together, these repositories represent the gold standard for time series classification research, with the UCR archive focusing on univariate sequences and the UEA archive addressing more complex real-world scenarios where multiple variables are tracked concurrently.
@@
For segmentation tasks, we employ the Child Mind Institute’s "Detect Sleep States" dataset [Esper et al., 2023], which consists of approximately 500 multi-day accelerometer recordings captured from wrist-worn devices, with each series corresponding to a unique participant. The dataset provides expert annotations for sleep onset and wakeup events, enabling comprehensive evaluation of temporal segmentation capabilities in physiological monitoring applications.
For anomaly detection, we incorporate several benchmark datasets representing different operational contexts and domains. These include NASA’s Mars Science Laboratory (MSL) and Soil Moisture Active Passive (SMAP) telemetry datasets [Hundman et al., 2018], which provide spacecraft sensor data with labeled anomalies; eBay’s Pooled Server Metrics (PSM) dataset [Abdulaal et al., 2021], containing server performance metrics; the Server Machine Dataset (SMD) [Su et al., 2019], which includes machine operation data; and the Secure Water Treatment (SWaT) dataset [Mathur and Tippenhauer, 2016], representing critical infrastructure monitoring. This diverse collection ensures robust anomaly detection capabilities across industrial, space, and cybersecurity applications.
@@
We also utilize the PTB-XL dataset [Wagner et al., 2022, Goldberger et al., 2000], which consists of 21,799 clinical 12-lead ECGs collected from 18,869 patients. Each recording spans 10 seconds and includes expert annotations from up to two cardiologists, using 71 distinct SCP-ECG standard statements. The dataset includes comprehensive metadata including patient demographics and diagnostic information, making it particularly valuable for developing medical time series analysis capabilities. From this dataset, we developed specialized variants optimized for different tasks: one for time series classification tasks and another for time series forecasting applications, as detailed in our data synthesis section.
@@
To ensure TsLLM maintains strong text generation capabilities alongside its time series analysis skills, we incorporate two standard text datasets into our training mixture. We utilize a subset of FineWeb-Edu [Lozhkov et al., 2024], a large-scale educational text corpus containing high-quality educational content curated from various web sources and specifically filtered for instructional content, explanations, and academic materials. Additionally, we include SmolTalk [Allal et al., 2024], a synthetic dataset of 1M instruction-following examples created for supervised fine-tuning, covering tasks like summarization and reasoning. This combination ensures that our model’s natural language capabilities remain robust while learning time series-specific skills.
C.3 Paired Time Series and Text Data
@@
Existing datasets that pair time series with textual data have been instrumental in developing models that can bridge numerical and linguistic understanding. The ECG-QA dataset [Oh et al., 2023] provides a specialized medical dataset that pairs electrocardiogram readings with natural language questions and answers about cardiac conditions and diagnoses. Built from PTB-XL [Wagner et al., 2022] and MIMIC-IV-ECG [Gow et al., 2023] databases, it includes both templated and naturally-phrased questions that test understanding of ECG features like heart rhythm patterns.
The TextTime Corpus (TTC) [Kim et al., 2024] offers broader domain coverage by pairing two types of temporal data: (1) daily weather measurements with forecast discussions and (2) patient vital signs paired with clinical notes from MIMIC-III [Johnson et al., 2016]. This diversity in both domains and text types provides valuable training examples for contextual reasoning about time series data.
Time-MMD [Liu et al., 2024c] extends multimodal capabilities by combining numerical and textual sequences across 9 domains, incorporating corresponding reports and search results that provide rich contextual information for time series analysis.
@@
We constructed this dataset using news data from two sources with different granularities: The Bloomberg dataset [Genloop, 2024] provides full articles with multiple stock symbol associations, while the Benzinga dataset [Aenlle, 2020] contributes headlines with single stock symbol associations. The dataset includes daily closing prices and company fundamentals (sector, industry, country, and business descriptions). Each forecasting task combines historical price data and news context (up to 4 Bloomberg articles or 16 Benzinga headlines) to predict a 32-day forward price sequence. We employ a minimum context window of 32 days with prediction windows spaced at 16-day intervals, ensuring all contextual information precedes the target sequence.
C.4 Synthetic Data Generation

### v2

@@ A.1 Classification
For classification tasks, we evaluate our model on the UCR-TSC dataset. Performance is assessed using the mean and median accuracy across sub-datasets, ensuring robustness. Results can be found in Table 6. TsLLM achieves competitive performance (81.2% mean accuracy with few-shot prompting) without task-specific architectures, demonstrating effective transfer of language model capabilities to temporal pattern recognition.
Table 6: Time series classification accuracy on the UCR-TSC benchmark collection. Although TsLLM is not designed for general numerical-only tasks, it performs comparably with top models.
Metric DTW CNN T-Loss TS2Vec Times- MOM- TsLLM TsLLM
@@
Mean 0.764 0.751 0.833 0.851 0.572 0.794 0.792 0.812
Median 0.768 0.773 0.849 0.871 0.565 0.815 0.798 0.823
A.2 Forecasting
We evaluate forecasting performance on the GIFT-Eval dataset. Performance is measured using a relative Mean Absolute Percent Error (rMAPE) normalized against the seasonal naive baseline. All baseline results are as reported in Aksu et al. [25]. We further break down performance by the number of variates (univariate vs multivariate) and prediction length, as illustrated in Table 7, where TsLLM outperforms competitively with other models although not completely outperforming all.
Table 7: Forecasting performance on GIFT-Eval benchmark across different data characteristics. Values show relative MAPE (lower is better) normalized to seasonal naive baseline, following [25]. TsLLM with few-shot prompting (FS) achieves best-in-class performance on univariate series (0.78) while remaining competitive overall (0.87), with particularly strong results on long-horizon forecasts.
@@ A.2 Forecasting
Anomaly Detection
@@
We also evaluate TsLLM on anomaly detection, specifically on the MSL dataset [35], as shown in Table 8. While our approach achieves a competitive F1 score of 73.3, there remains a performance gap compared to state-of-the-art models like UniTS-FT (81.2) and iTransformer-FT (80.4). The results demonstrate TsLLM’s versatility across different time series tasks, though specialized models may still hold advantages in specific domains.
@@ Appendix B Datasets
B.1 Training Data Overview
Table 9 summarizes the key statistics for all datasets used in training TsLLM. Our training corpus encompasses nearly 27 billion total tokens across 20 distinct datasets, spanning diverse domains including medical signals, financial data, environmental measurements, and industrial sensors. The datasets support multiple task types including forecasting, classification, anomaly detection, segmentation, and question-answering, enabling TsLLM to develop comprehensive time series analysis capabilities.
@@ B.1 Training Data Overview
GIFT-Pretrain Multi-Domain Forecasting 7,951,924,108 3,861,122,236 (1,990,801,872/5,961,122,236) 74,805,850
ECG-QA ECG Q-A 1,071,939,892 10,119,218 (468,584,812/603,355,080) 942,950
@@
PTBXL-Classification ECG Classification 20,615,930 296,710 (3,577,588/16,741,632) 21,799
PTBXL-Forecasting ECG Forecasting 22,482,365 5,740,733 (0/16,741,632) 21,799
PTBXL-Report Gen ECG Report Gen. 23,155,582 3,689,075 (2,724,875/16,741,632) 21,799
Monster Multi-Domain Classification 1,665,431,147 12,045,400 (929,008,270/736,422,877) 1,998,336
@@
Synthetic Multi-Domain Mixed 15,186,959,206 2,815,439,507 (11,267,019,855/3,919,939,351) 11,273,306
Total All All 26,935,940,241 7,287,761,930 (15,521,996,808/11,340,414,501) 94,062,682
@@
B.2 Unimodal Time Series Data
@@
For large-scale pretraining, we utilize the GIFT-Eval benchmark [25], a comprehensive collection of 144,000 time series across 23 datasets spanning seven domains. GIFT-Eval provides dedicated splits for pretraining, training, and testing, making it particularly suitable for developing and evaluating general-purpose time series models. The benchmark’s temporal granularity ranges from secondly to yearly frequencies, incorporating both univariate and multivariate sequences with forecast horizons varying from short-term to long-term predictions. With over 230 billion data points, this dataset provides the scale and diversity necessary for training foundation models. We use it extensively for pre-training our time series encoder and decoder models, as well as training TsLLM for forecasting tasks, while carefully ensuring no data leakage into the GIFT-Eval test set used in our evaluation (Section 5).
@@
For time series classification tasks, we employ two principal benchmark collections that have become standard in the field. The UCR Time Series Classification Archive [36] comprises 128 datasets spanning diverse applications including electrocardiograms, motion capture sequences, sensor readings, and spectrographs. These datasets vary significantly in length, class distribution, and domain-specific characteristics, providing comprehensive coverage for univariate time series classification. Complementing this, the UEA Time Series Classification Archive [37] extends our capabilities to multivariate time series, offering 30 datasets that capture complex temporal patterns in applications such as human activity recognition, motion capture, and handwriting analysis. Together, these repositories represent the gold standard for time series classification research, with the UCR archive focusing on univariate sequences and the UEA archive addressing more complex real-world scenarios where multiple variables are tracked concurrently.
@@
For segmentation tasks, we employ the Child Mind Institute’s "Detect Sleep States" dataset [38], which consists of approximately 500 multi-day accelerometer recordings captured from wrist-worn devices, with each series corresponding to a unique participant. The dataset provides expert annotations for sleep onset and wakeup events, enabling comprehensive evaluation of temporal segmentation capabilities in physiological monitoring applications.
For anomaly detection, we incorporate several benchmark datasets representing different operational contexts and domains. These include NASA’s Mars Science Laboratory (MSL) and Soil Moisture Active Passive (SMAP) telemetry datasets [35], which provide spacecraft sensor data with labeled anomalies; eBay’s Pooled Server Metrics (PSM) dataset [39], containing server performance metrics; the Server Machine Dataset (SMD) [40], which includes machine operation data; and the Secure Water Treatment (SWaT) dataset [41], representing critical infrastructure monitoring. This diverse collection ensures robust anomaly detection capabilities across industrial, space, and cybersecurity applications.
@@
We also utilize the PTB-XL dataset [42, 43], which consists of 21,799 clinical 12-lead ECGs collected from 18,869 patients. Each recording spans 10 seconds and includes expert annotations from up to two cardiologists, using 71 distinct SCP-ECG standard statements. The dataset includes comprehensive metadata including patient demographics and diagnostic information, making it particularly valuable for developing medical time series analysis capabilities. From this dataset, we developed specialized variants optimized for different tasks: one for time series classification tasks and another for time series forecasting applications, as detailed in our data synthesis section.
@@
To ensure TsLLM maintains strong text generation capabilities alongside its time series analysis skills, we incorporate two standard text datasets into our training mixture. We utilize a subset of FineWeb-Edu [44], a large-scale educational text corpus containing high-quality educational content curated from various web sources and specifically filtered for instructional content, explanations, and academic materials. Additionally, we include SmolTalk [28], a synthetic dataset of 1M instruction-following examples created for supervised fine-tuning, covering tasks like summarization and reasoning. This combination ensures that our model’s natural language capabilities remain robust while learning time series-specific skills.
B.3 Paired Time Series and Text Data
@@
Existing datasets that pair time series with textual data have been instrumental in developing models that can bridge numerical and linguistic understanding. The ECG-QA dataset [29] provides a specialized medical dataset that pairs electrocardiogram readings with natural language questions and answers about cardiac conditions and diagnoses. Built from PTB-XL [42] and MIMIC-IV-ECG [45] databases, it includes both templated and naturally-phrased questions that test understanding of ECG features like heart rhythm patterns.
The TextTime Corpus (TTC) [46] offers broader domain coverage by pairing two types of temporal data: (1) daily weather measurements with forecast discussions and (2) patient vital signs paired with clinical notes from MIMIC-III [47]. This diversity in both domains and text types provides valuable training examples for contextual reasoning about time series data.
Time-MMD [48] extends multimodal capabilities by combining numerical and textual sequences across 9 domains, incorporating corresponding reports and search results that provide rich contextual information for time series analysis.
@@
We constructed this dataset using news data from two sources with different granularities: The Bloomberg dataset [49] provides full articles with multiple stock symbol associations, while the Benzinga dataset [50] contributes headlines with single stock symbol associations. The dataset includes daily closing prices and company fundamentals (sector, industry, country, and business descriptions). Each forecasting task combines historical price data and news context (up to 4 Bloomberg articles or 16 Benzinga headlines) to predict a 32-day forward price sequence. We employ a minimum context window of 32 days with prediction windows spaced at 16-day intervals, ensuring all contextual information precedes the target sequence.
B.4 Synthetic Data Generation